jovanie:

the legs of a right triangle are 5 units and 15 units. what is the length of the hypotenuse

Mercury:

pythagorean theorem a^2 + b^2 = c^2 where a and b are legs, and c is the hypotenuse in your case, a = 5 and b = 15 and you can plug these in to solve for c remember, the formula gives you c^2 not just c, so you'll have to take the square root at the end

jovanie:

please help me

jovanie:

so whats the answer

Mercury:

please be mindful we are a teaching site, and we don't give direct answers here please re-read my comment, and attempt to solve. if you are still confused, please notify me and I'll try to simplify further

foxy44:

ya what she said

foxy44:

or he

jovanie:

bruh i need help please this stuff makes now sense

Mercury:

a^2 + b^2 = c^2 where a and b are legs, and c is the hypotenuse since your legs are 5 and 15, you can plug these in as your a and b values to get: 5^2 + 15^2 = c^2 can you try adding 5^2 and 15^2 together?

foxy44:

I would say but the only way you succeed is if you try and that's how your brain grows.

Mercury:

5^2 = 5*5 15^2 = 15*15 so 5*5 + 15*15 = ?

jovanie:

40

jovanie:

mecury

Mercury:

not quite. remember * means multiply 5*5 = 25 15*15 = 225 25 + 225 = ?

jovanie:

250

foxy44:

yes 250

jovanie:

there is two 250

Mercury:

good, so that's your c^2 value since c^2 = 250, take the square root of both sides to get c what's the square root of 250? use a calculator if you are unsure

jovanie:

15

jovanie:

mecury

Mercury:

sqrt(250) is about 15.8, yes

Mercury:

does it tell you what format the answer should be in? like, decimal or radical form?

jovanie:

it says 250 and sqaure root 250

Mercury:

then sqrt(250) should be your solution

jovanie:

mecury